Alerts have now resolved. Moving back to fully operational status.
Nov 29, 14:22 MST
Service has now been restored. Our team is investigating to verify stability.
Nov 29, 14:15 MST
The data centre network team believes they have located the issue and have notified us that network service should be restoring soon.
Nov 29, 14:06 MST
Outage still in effect. Network team at the data centre verified that the switches were rebooted, but there are additional issues affecting the network that the network team is investigating.
Nov 29, 13:52 MST
Outage still in effect. Network team at the data centre is working to resolve the issue.
Nov 29, 13:36 MST
Outage has now lasted ~20 minutes -- 3:00 PM EDT to 3:20 PM EDT. We are engaging with the East data centre operational staff to determine the situation.
Nov 29, 13:22 MST
Our east data centre has just notified us of emergency switch maintenance from 3:00 PM EDT to 4:00 PM EDT. Switch reboots during this time may cause outages of up to 5 minutes at a time.
Nov 29, 13:08 MST